OpenStudy (anonymous):

log2(a^2-6a)=log2(10+3a)

OpenStudy (lgbasallote):

you can cancel out the log_2 if i'm not mistaken since both sides have them

OpenStudy (anonymous):

\[Log2(a ^{2}-6a)=Log2(10+3a)\] \[Log2/Log2=(10+3a)/(a ^{2}-6a)\] \[1=(10+3a)/(a ^{2}-6a)\] \[a ^{2}-6a=10+3a\] \[a ^{2}-6a-3a-10=0\] \[a ^{2}-9a-10=0\] \[a ^{2}+a-10a-10=0\] \[a(a+1)-10(a+1)=0\] \[(a+1)(a-10)=0\] \[a+1=0\] \[a=-1\] Or, \[a-10=0\] \[a=10\] Therefore, 10 and -1 are the roots of the solution.

OpenStudy (mertsj):

if the bases are the same and the exponents (logs) are the same, then the numbers must be the same number. So: \[a^2-6a=10+3a\] And \[a^2-9a-10=0\] Then proceed as Vengeance has shown